### v4

# LLMLight: Large Language Models as Traffic Signal Control Agents

Abstract.
Traffic Signal Control (TSC) is a crucial component in urban traffic management, aiming to optimize road network efficiency and reduce congestion. Traditional methods in TSC, primarily based on transportation engineering and reinforcement learning (RL), often exhibit limitations in generalization across varied traffic scenarios and lack interpretability. This paper presents LLMLight, a novel framework employing Large Language Models (LLMs) as decision-making agents for TSC. Specifically, the framework begins by instructing the LLM with a knowledgeable prompt detailing real-time traffic conditions. Leveraging the advanced generalization capabilities of LLMs, LLMLight engages a reasoning and decision-making process akin to human intuition for effective traffic control. Moreover, we build LightGPT, a specialized backbone LLM tailored for TSC tasks. By learning nuanced traffic patterns and control strategies, LightGPT enhances the LLMLight framework cost-effectively. Extensive experiments on nine real-world and synthetic datasets showcase the remarkable effectiveness, generalization ability, and interpretability of LLMLight against nine transportation-based and RL-based baselines.
traffic signal control, large language model, traffic control agent, intelligent transportation
$*{}^{*}$ Corresponding author.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Planning for deterministic actions
Figure 1. The workflow of LLMLight.
1. Introduction
Traffic congestion has emerged as a critical issue impacting both human society and the environment. As urban migration continues to accelerate, city populations are expanding rapidly, further intensifying this challenge. This trend has necessitated a focused effort on optimizing Traffic Signal Control (TSC), a pivotal research area in intelligent transportation (Wu et al., 2023; Zhang et al., 2023). Efficient TSC not only promises substantial economic and environmental benefits but also enhances societal well-being. However, effectively managing traffic signals in complex and dynamic urban road networks remains a daunting task.
Existing research on TSC has primarily fallen into two categories: transportation (Koonce and Rodegerdts, 2008; Hunt et al., 1982; Lowrie, 1990; Varaiya, 2013; Cools et al., 2013) and Reinforcement Learning (RL)-based approaches (Chen et al., 2020; Wei et al., 2019b; Oroojlooy et al., 2020; Chenguang et al., 2021; Wu et al., 2021; Zhang et al., 2022b; Wei et al., 2019a; Wu et al., 2023). Transportation methods primarily focus on crafting efficient heuristic algorithms, dynamically adapting traffic signal configurations based on lane-level traffic conditions. However, these methods heavily rely on manual design, demanding substantial human effort. The emergence of deep neural networks (DNNs) (Mnih et al., 2015) led to the introduction of deep RL-based techniques to address this challenge (Abdulhai et al., 2003; El-Tantawy and Abdulhai, 2010). These approaches have exhibited remarkable performance across various traffic scenarios. Nevertheless, RL-based methods also present several drawbacks. Primarily, they may struggle with limited generalization ability, particularly when transferring to larger-scale road networks or under highly uncommon scenarios (e.g., extreme high-traffic situations), as their training data only covers limited traffic situations. Additionally, RL-based methods lack interpretability due to the black-box nature of DNNs, which makes it hard to explain the rationale behind their control actions under specific traffic conditions.
The recent emergence of Large Language Models (LLMs) has revolutionized several domains with their exceptional zero-shot learning and generalization capabilities, emulating human-like reasoning to solve complex tasks. For instance, AutoGPT (et al., 2023) proposes to break tasks into multiple sub-goals and iterate until the main task is completed. In the transportation domain, GPT-Driver (Mao et al., 2023) proposes to instruct GPT-3.5 to tackle motion planning tasks in autonomous driving. For traffic control, PromptGAT (Da et al., 2023) uses LLMs to generate human knowledge to help the DNN model understand long-tail cases (e.g., extreme weather) in TSC tasks, aiming to bridge the gap between real-world and simulations. However, while existing research utilizes LLMs as auxiliary tools to enhance decision-making, the direct employment of LLMs as TSC agents for human-like decision-making remains largely unexplored. In particular, two major challenges arise toward this goal.
First, LLMs are typically pre-trained on large-scale natural language corpora and rarely incorporate non-textual traffic data, such as sensor readings and GPS trajectories (Touvron et al., 2023; Chung et al., 2022). Despite their generalization capability across various tasks and domains, an inherent gap exists between real-time traffic data and linguistic understanding. The first challenge lies in enabling LLMs to comprehend real-time traffic dynamics and effectively interact with the traffic environment, which is critical for effective LLM-based traffic signal control.
Second, selecting and developing an effective LLM for TSC poses another significant challenge. Generalist LLMs often lack specific domain knowledge and are prone to hallucination problems in professional fields (Huang et al., 2023; Ji et al., 2023). Although state-of-the-art LLMs such as GPT-4 demonstrate promising generalization abilities, their closed-source nature and substantial usage costs pose barriers to their optimization for real-time TSC tasks. Consequently, building a specialized LLM tailored for TSC tasks is crucial to deliver more effective and human-aligned control policies.
To this end, we introduce LLMLight, a traffic signal control agent framework based on LLMs, as depicted in Figure 1. Specifically, we consider TSC as a partially observable Markov Game, where each agent, armed with an LLM, manages the traffic light at an intersection. At each signal-switching time step, the agent collects traffic conditions of the target intersection and transforms them into human-readable text as real-time observation. Additionally, we incorporate task descriptions enriched with commonsense knowledge about a control strategy to aid the LLM’s understanding of traffic management tasks. The combined real-time observation, task description, and control action space form a knowledgeable prompt that guides the LLM’s decision-making. Then, the agent leverages Chain-of-Thought (CoT) reasoning to determine the optimal traffic signal configuration for the subsequent time step.
Furthermore, we construct a specialized LLM, LightGPT, to enhance the LLMLight framework. On the one hand, we propose imitation fine-tuning to let the specialized LLM exploit high-quality control actions and underlying rationales derived from GPT-4. On the other hand, we introduce a policy refinement process that utilizes a well-trained critic model to evaluate and improve the agent’s actions. The optimized LightGPT produces more effective control policies and showcases remarkable generalization ability across diverse traffic scenarios in a more cost-effective manner than GPT-4.
We summarize the major contributions of this paper as follows: (1) We pioneer the use of advanced LLMs as control agents, empowering human-like decision-making for interpretable traffic signal control. To our knowledge, this is the first exploration of LLMs as decision-making agents in TSC. (2) We introduce LLMLight [^1] Our project is available at https://github.com/usail-hkust/LLMTSCS, a tailored framework that engages various LLMs in effectively regulating traffic signals via knowledgeable instructions. (3) We construct LightGPT [^2] The model weights is available at https://huggingface.co/USAIL-HKUSTGZ/LLMLight-LightGPT, a specialized LLM backbone, which is optimized for TSC tasks and significantly enhances the framework’s capability to derive effective control policies in a cost-effective manner. (4) Extensive experiments on nine traffic flow datasets demonstrate the superior effectiveness, generalization capabilities, and interpretability of our framework. Additionally, these experiments reveal key insights into the potential of integrating LLMs into intelligent transportation systems.
2. Preliminary
In this section, we first introduce the key concepts in traffic signal control tasks.
Definition 2.1 ().
Road network: The road network is a directed graph consisting of intersection joints $I$ and lanes $L$. Lanes are categorized into go-through ($L_{go}$), left-turn ($L_{left}$), and right-turn ($L_{right}$) lanes, determined by the movement of vehicles on the lane. Each lane is divided into multiple segments $S=\left\{s_{1},\dots,s_{n}\right\}$.
Definition 2.2 ().
Traffic signal control action: The traffic signal control action corresponds to a specific signal phase, denoted as $a=set(L_{allow})$, where $L_{allow}$ is a group of allowed lanes without conflicting movements. $set(L_{allow})$ indicates the green light for activating the signal phase. At each signal-switching time step, the agent controlling the intersection chooses an action from $A=\left\{a_{1},\dots,a_{m}\right\}$. The active signal phase allows vehicles in a specific lane group to pass, while other phases display the red light.
Please refer to Appendix A.1 (Figure 9) for more details of the most used setting of the intersection, lanes, and signal phases.
3. LLM for Traffic Signal Control
This section first introduces the task formulation of traffic signal control in the LLM-empowered context. Then, we delve into the workflow of our proposed LLMLight, elucidating the methodology for prompting LLMs to perceive, analyze, and regulate traffic at the target intersection directly. Finally, we elaborate on the training procedures implemented for developing LightGPT, aiming to augment the backbone LLM’s domain expertise in traffic management, thereby generating proficient control policies adaptable to diverse traffic scenarios.
3.1. Task Formulation
We define traffic signal control as a partially observable Markov Game (Wei et al., 2019b, a). The LLM-based agent manages the traffic light of an intersection with policy $\pi_{\theta}$. Based on the observation $o$ of the real-time traffic condition at the intersection, traffic scenario description $d_{scene}$, task descriptions $d_{task}$, commonsense knowledge $d_{know}$, and control action space $A$, the LLM outputs the reasoning trajectory $Y$ with the control action $a$ that regulates the traffic signal of the intersection, with the long-term goals of maximizing the transportation efficiency within the road network:
(1) $(Y,a)=\pi_{\theta}(Prompt(o,d_{scene},d_{task},d_{know},A)).$
3.2. Workflow of LLMLight
This subsection details the workflow of our proposed LLMLight, including (1) observation collection: it collects the traffic condition of the intersection; (2) commonsense augmented prompt construction: it composes a commonsense knowledge enhanced prompt, instructing the LLM to devise the optimal traffic signal configuration; (3) reasoning and action execution: the LLM engages a decision-making process by employing the constructed prompt, and subsequently executes its selected control action.
3.2.1. Observation Feature Construction
We collect traffic condition features that can be easily obtained from the real-world traffic control environment:
  • Queuing vehicle count: Vehicles with speeds slower than the threshold $v_{stop}$ are considered as queuing vehicles. We count their number in lane $l$ as $n^{l}_{q}$.
  • Approaching vehicle count: Vehicles faster than $v_{stop}$ are considered as approaching vehicles. We count the number of approaching vehicles in the segment $s$ of lane $l$ as $n^{l}_{s}$.
Subsequently, these features on every incoming lane (where the vehicle enters the intersection) are transformed into human-readable text as the observation $o$ of the LLM.
(a) The prompt template.
(b) The reasoning and action execution process.
Figure 2. The prompt template of LLMLight and decision-making process of the backbone LLM. In this scenario, the intersection connects four lane sections, including north, south, east, and west.
3.2.2. Commonsense Augmented Prompt Construction
In addition to observation, we provide the LLM with comprehensive information essential in addressing traffic signal control, including traffic scenario description $d_{scene}$, task descriptions $d_{task}$, and control action space $A$. This enables the LLM to thoroughly understand the task, facilitating reasonable control actions. Furthermore, we integrate commonsense knowledge $d_{know}$ to mitigate generalist LLMs’ limitations in traffic management, a domain they are not explicitly trained for. Specifically, this knowledge dictates that the agent needs to prioritize lanes with extended queue lengths while paying less attention to vehicles distant from the intersection. Formally, we denote the composed prompt as $X=Prompt(o,d_{scene},d_{task},d_{know},A)$. A glimpse of the prompt template is shown in Figure 1(a), and a detailed example is provided in Appendix A.5 (Table 4).
3.2.3. Reasoning and Action Execution
We instruct the LLM to perform zero-shot reasoning using the constructed prompt. Its decision-making unfolds in two key steps: reasoning and action execution. Initially, it embarks on comprehensive analysis, assimilating the given task, leveraging commonsense knowledge, and assessing the current traffic conditions across various lanes. Subsequently, it selects the signal for the lanes experiencing the highest congestion, thereby optimizing traffic flow and ensuring the smooth passage of vehicles. In this manner, LLMLight can not only make effective control actions but also distinguishes itself by providing detailed rationales for each decision, contributing to a more explainable and transparent traffic control system. Formally, we denote the reasoning and action execution as $(Y,a)=\pi_{\theta}(X)$. An example of the backbone LLM’s decision-making process is shown in Figure 1(b). We conduct an interpretability analysis in Subsection 4.7.
3.3. Backbone LLM Training
Fine-tuning has become a common routine to align LLM with diverse downstream applications (Chen et al., 2023a; Zeng et al., 2023) by augmenting the domain expertise in the specific realm. We propose a training procedure to optimize an LLM specifically tailored for traffic signal control, leading to the development of LightGPT. Figure 3 presents the training workflow. It primarily consists of three stages: (1) reasoning trajectory collection and filtering: it collects the CoT reasoning trajectory of GPT-4 for imitation fine-tuning, then we select the trajectories that most align with the long-term goal to ensure the data quality; (2) imitation fine-tuning: it leverages the collected reasoning trajectories for training the backbone LLM on actions taken by GPT-4 and their underlying rationales; (3) critic-guided policy refinement: it further improves the decision-making process of the LLM by fine-tuning with the feedback of a well-trained critic model.
Figure 3. The training procedure of LightGPT.
3.3.1. Trajectory Collection and Filtering
Following the prompt construction process detailed in Section 3.2.2, we first instruct GPT-4 to interact with simulated traffic environments for gathering reasoning trajectories for subsequent imitation fine-tuning. To ensure the quality of the collected data, we employ a filtering strategy that identifies reasoning trajectories aligning most closely with the long-term objective of traffic signal control, such as minimizing future queue length. This filter operation is achieved by aligning with a pre-trained action-value network (the architecture is derived from (Zhang et al., 2022b)), which is trained by optimizing Bellman equation within the simulation environment:
(2) $\mathcal{L}_{value}=\mathbb{E}[(R(o_{t},a_{t})+\gamma\max\limits_{a_{t+1}}Q(o_{t+1},a_{t+1})-Q(o_{t},a_{t}))^{2}],$
where $o_{t}$ and $a_{t}$ are the observation and action chosen at the signal-switching time step $t$, $\gamma\in[0,1]$ is the discount factor. $R(o,a)$ is the reward function that provides feedback for executing action $a$ under observation $o$ (e.g., the negative of queue length). $Q(o,a)$ is the action-value function estimating the cumulative future reward earned after executing $a$. Actions estimated to yield higher future rewards suggest greater alignment with the long-term goal.
Subsequently, the trained action-value function is utilized as a critic to evaluate the decision-making of GPT-4. We selectively retain only reasoning trajectories where the selected actions are associated with the highest estimated future reward, as:
(3) $\big{\{}(X_{t},Y_{t},a_{t})|a_{t}=\mathop{argmax}\limits_{a\in A}Q(o_{t},a)\big{\}}^{T}_{t=0},$
where $T$ is the simulation duration, $X_{t}$ is prompt, and $Y_{t}$ is the reasoning trajectory of GPT-4.
3.3.2. Imitation Fine-tuning
In the initial stage of backbone LLM training, we employ an imitation learning process, allowing the LLM to train on the control actions taken by GPT-4 and their underlying rationales. We treat prompt $X$ as fine-tuning instructions and GPT-4’s reasoning trajectory $Y=[Y;a]$ containing its selected action $a$ as the anticipated response. Negative log-likelihood (NLL) is adopted as the loss function:
(4) $\mathcal{L}_{IFT}=-\sum_{w}^{|Y|}logP_{\pi_{\theta}}(y_{w}|X,Y_{<w}),$
where $P_{\pi_{\theta}}(y_{w}|X,Y_{<w})$ is the probability of being prompted with $X$ to generate the token $y_{w}$. Due to the enormous GPU memory usage of full model training, we adopt parameter-efficient fine-tuning (PEFT) techniques, specifically low-rank adaptation (LoRA) (Hu et al., 2022), in our training. LoRA only adds small adapters on top of some specific layers, drastically reducing the number of trainable parameters.
3.3.3. Critic-guided Policy Refinement
To further enhance the effectiveness of the LLM’s control policy, we propose a policy refinement procedure, which adjusts the LLM’s reasoning trajectory for directing better actions. Similarly, we use the pre-trained action-value function elaborated in Section 3.3.1 as the critic model to evaluate the actions derived by the backbone LLM. Higher evaluation scores correspond to greater estimated future rewards. Subsequently, we implement an alignment fine-tuning algorithm to adjust the reasoning trajectory, ultimately guiding the LLM towards actions that yield higher scores. Inspired by studies like (Wang et al., 2023a; Yuan et al., 2023), we leverage a ranking loss function to achieve this goal.
Specifically, we have $k$ different reasoning trajectories $\mathbf{Y}=\left\{Y_{i}\right\}^{k}_{i=1}$ sampled by policy $\pi_{\theta}$ under instruction $X$. The critic gives scores for actions derived by each trajectory $Y_{i}$ as $q_{i}=Q(o,a_{i})$. Afterwards, the token-averaged log-likelihood value of each $Y_{i}$ is derived to denote the corresponding probability for $\pi_{\theta}$ to generate $Y_{i}$:
(5) $p_{i}=\frac{1}{{|Y_{i}|}}\sum^{|Y_{i}|}_{w}logP_{\pi_{\theta}}(y_{i,w}|X,Y_{i,<w}).$
Following (Wang et al., 2023a), we adopt ranking-feedback loss with a boundary-constrained term (RBC) for generating reasoning trajectories that result in actions with higher scores:
(6) $\mathcal{L}_{CGPR}=log\big{\{}1+\sum_{q_{i}>q_{j}}\big{[}e^{(p_{j}-p_{i})}+e^{(2p_{j*}-2\beta-p_{i}-p_{j})}\big{]}\big{\}},$
where $p_{j*}=\min\limits_{q_{k}>q_{j}}p_{k}$ is the probability of the least favorable reasoning trajectory rated higher than $Y_{j}$, $\beta$ is a hyper-parameter, $e^{(p_{j}-p_{i})}$ is the alignment term to elevate the probability of generating trajectories associated with actions rated higher scores, and $e^{(2p_{j*}-2\beta-p_{i}-p_{j})}$ is the constraint term for prohibiting performance degradation.
It’s crucial to aware that while this refinement procedure is guided by a pre-trained critic model developed from RL, this doesn’t imply that the backbone LLM learns exclusively from it. The alignment process is trained on reasoning trajectories sampled by the LLM’s policy. The objective is to encourage the model to discern which trajectory is more reasonable and leads to decisions more closely aligned with long-term goals. However, it’s important to note that we do not provide any ground-truth labels, thus the LLM is not bound to follow the critic’s predicted “best” action directly. The experiments presented in Subsection 4.6 validate that RL models experience performance degradation under varied traffic scenarios, whereas LLMLight equipped with LightGPT remains robust based on our proposed training procedure.
4. Experiments
In this section, we conduct extensive experiments to evaluate our proposed LLMLight by answering the following research questions:
  • RQ1: How is the effectiveness of LLMLight compared with traditional transportation and RL methods?
  • RQ2: How is the generalization ability of LLMLight across different traffic scenarios and vehicle volumes?
  • RQ3: How is the interpretability of LLMLight in providing rationales behind the selected control actions?
4.1. Experimental Settings
Table 1. Statistics of datasets.
Flow dataset # of inter- # of vehicles Arrival rate (vehicles/5min)
sections Mean Std Max Min
Jinan 1 12 6295 523.67 98.52 671 255
Jinan 2 4365 362.83 74.81 493 236
Jinan 3 5494 456.92 46.38 544 362
Jinan Extreme 24000 1999.08 44.14 2097 1932
Hangzhou 1 16 2983 247.67 40.44 332 211
Hangzhou 2 6984 581.08 318.43 1145 202
Hangzhou Extreme 32000 2665.75 46.89 2722 2550
New York 1 196 11058 849.69 174.06 964 382
New York 2 16337 1255.77 264.85 1440 475
4.1.1. Environment Settings
We conduct experiments on an open-sourced simulator, CityFlow (Zhang et al., 2019), to evaluate the efficiency of compared methods. It provides extensive APIs to access the traffic state features and execute the selected control actions. In our experiments, each intersection connects four lane sections: East, West, North, and South. We utilize four signal phases as corresponding control actions, including ETWT (go-through from east and west), ELWL (left-turn from east and west), NTST (go-through from north and south), and NLSL (left-turn from north and south). The turn-right movements are always allowed in simulation. The green signal phase duration is set to thirty seconds. Following existing studies (Zhang et al., 2022b; Wei et al., 2019b), a three-second yellow signal and a two-second all-red time follows each green signal to prepare the transition. The simulation spans one hour.
4.1.2. Datasets
Our experiments use seven real-world (Wei et al., 2019c) and two synthetic traffic flow datasets. Statistics for these datasets are detailed in Table 1. Visualizations of road networks are presented in Appendix A.4 (Figure 11).
  • Jinan: These datasets are collected from the Dongfeng sub-district, Jinan, China, featuring 12 intersections. The East and West sections span 400 meters, while the South and North sections cover 800 meters. Across different periods, there are three traffic flow datasets recorded in this area.
  • Hangzhou: These datasets are collected in the Gudang sub-district, Hangzhou, China. There is a total of 16 intersections in this area. The East and West sections span 800 meters, while the South and North sections extend 600 meters. Across different periods, this dataset comprises two traffic flow datasets from the recorded area.
  • New York: To evaluate the scalability of different methods, we introduce two large-scale traffic datasets collected in the Upper East Side of Manhattan, New York, USA, utilizing publicly available taxi trip data. It encompasses 196 intersections within the region. All lanes span 300 meters. Over different periods, the dataset includes two distinct traffic flow datasets.
  • Extreme High-traffic Sceanrios: To simulate high-traffic scenarios, our experiments encompass two synthetic traffic flow datasets on the Jinan and Hangzhou road networks, featuring about four times more vehicles arriving within each 5-minute interval than the original datasets.
4.2. Metrics
Following previous studies (Zhang et al., 2022b), we leverage average travel time (ATT), average queue length (AQL), and the average waiting time (AWT) of vehicles to evaluate the performance of different policies made by traffic signal control agents.
  • Average traveling time (ATT): The average traveling time quantifies the average duration of all the vehicles traveling from their origins to their respective destinations.
  • Average queue length (AQL): The average queue length is defined as the average number of queuing vehicles waiting in the road network.
  • Average waiting time (AWT): The average waiting time quantifies the average queuing time of vehicles at every intersection in the road network.
4.3. Compared Models
For transportation methods, we adopt Random, FixedTime, and Maxpressure as baselines. For RL methods, we compare MPLight, AttendLight, PressLight, CoLight, Efficient-CoLight, and Advanced-CoLight with our proposed method. For LLMs, we leverage GPT-4, ChatGPT-3.5, and the series of Llama-2 models as the backbone LLM (without fine-tuning) of LLMLight. Please refer to Appendix A.2 for more details about the abovementioned models.
4.4. Model Settings
All RL methods are trained with uniform hyperparameters, encompassing a learning rate of $1\times 10^{-3}$, a replay buffer size of $12000$, a sample size of $3000$, and a hidden size of $20$. The top_p value is standardized at $1.0$ for all LLMs. To stabilize the LLMs’ output, a temperature setting of zero is implemented for ChatGPT-3.5 and GPT-4, while Llama-2 is set to $0.1$. The LightGPT is developed from Llama-2-13B. In the imitation fine-tuning, the batch and mini-batch sizes are set to $128$ and $16$, respectively. We sample four reasoning trajectories under each instruction for critic-guided policy refinement. The fine-tuning learning rate is set to $3\times 10^{-4}$.
Table 2. Overall performances of LLMLight and previous traditional methods on Jinan and Hangzhou datasets. The best, second-best, and third-best results are highlighted through boldface, double underline, and underline, respectively.
Method Jinan Hangzhou
1 2 3 1 2
ATT AQL AWT ATT AQL AWT ATT AQL AWT ATT AQL AWT ATT AQL AWT
Transportation Methods
Random 597.62 687.35 99.46 555.23 428.38 100.40 552.74 529.63 99.33 621.14 295.81 96.06 504.28 432.92 92.61
FixedTime 481.79 491.03 70.99 441.19 294.14 66.72 450.11 394.34 69.19 616.02 301.33 73.99 486.72 425.15 72.80
Maxpressure 281.58 170.71 44.53 273.20 106.58 38.25 265.75 133.90 40.20 325.33 68.99 49.60 347.74 215.53 70.58
RL Methods
MPLight 307.82 215.93 97.88 304.51 142.25 90.91 291.79 171.70 89.93 345.60 84.70 81.97 358.56 237.17 100.16
AttendLight 291.29 186.25 61.73 280.94 115.52 52.46 273.02 144.05 55.93 322.94 66.96 55.19 358.81 239.05 72.88
PressLight 291.57 185.46 50.53 281.46 115.99 47.27 275.85 148.18 54.81 364.13 98.67 90.33 417.01 349.25 150.46
CoLight 279.60 168.53 58.87 274.77 108.28 54.14 266.39 135.08 53.33 322.85 66.94 61.82 342.90 212.09 99.74
Efficient-CoLight 277.11 163.60 43.41 269.24 102.98 39.74 262.25 129.72 39.99 311.96 58.20 36.83 333.27 189.65 61.70
Advanced-CoLight 274.67 160.85 49.30 268.25 102.12 41.11 260.66 127.83 43.54 304.47 52.94 41.75 329.16 186.34 76.59
LLMLight
Llama-2-13B 508.72 567.61 169.67 435.17 290.61 156.30 452.24 416.38 240.99 427.37 152.18 157.47 432.21 388.59 220.67
Llama-2-70B 353.03 286.52 85.96 324.52 162.34 99.87 320.41 210.13 100.90 357.95 93.21 106.63 361.53 250.69 121.58
ChatGPT-3.5 536.79 952.93 155.62 524.81 393.21 179.34 501.36 479.50 154.19 463.04 181.95 191.87 418.75 336.47 130.56
GPT-4 275.26 160.93 46.61 271.34 105.22 47.55 264.70 132.53 46.16 318.71 62.84 58.09 335.81 193.32 66.02
LightGPT (IFT) 277.60 166.17 48.21 272.24 104.93 48.78 261.34 129.27 45.02 316.4 61.55 49.92 333.42 190.43 68.99
LightGPT 274.03 159.39 43.24 266.94 100.46 40.34 260.17 127.08 41.00 311.76 57.89 42.84 330.71 189.09 64.16
4.5. Comparison Between the Traditional Methods and LLMLight (RQ1)
We first conduct experiments on Jinan and Hangzhou with five real-world traffic flow datasets. The comparison between traditional approaches and LLMLight is shown in Table 2. LightGPT (IFT) refers to the model trained exclusively through imitation fine-tuning. The experiment results reveal that LLMLight equipped with LightGPT consistently achieves state-of-the-art (SOTA) or comparable performance against all baselines. In detail, almost all transportation methods lag behind RL and LLMLight (LightGPT). Maxpressure demonstrates favorable performance in average waiting time (AWT) since it considers queuing vehicles the only privilege. The CoLight models exhibit the best performance among all RL methods, as they introduce neighboring information besides the target intersection. Although Advanced-CoLight, the SOTA RL method, outperforms LLMLight (LightGPT) on the Hangzhou datasets, it requires communication among neighboring intersections. On the contrary, LLMLight (LightGPT) demonstrates competitive results by leveraging observation features solely from the target intersection, underscoring its remarkable effectiveness.
As for LLMLight powered by generalist LLMs, we observe that GPT-4 outperforms all other LLMs and demonstrates competitive results against the most advanced RL methods, underscoring the effectiveness of our proposed prompting framework. Llama2-70B and 13B secure the second and third positions, indicating the adherence of LLMs to the scaling law (Kaplan et al., 2020) in the traffic signal control tasks. Surprisingly, ChatGPT-3.5 exhibits the least favorable performance. These experiment results indicate that generalist LLMs may lack expertise in traffic management, and there is a need to develop such domain-specific LLMs. The detailed analysis of their failure is elaborated in Appendix A.3.
By fine-tuning to imitate the reasoning trajectories of GPT-4, LightGPT (IFT) demonstrates comparable or even better performance than GPT-4. This highlights the innate capability of the backbone model, Llama-2-13B, to excel in traffic signal control tasks if not constrained by its absence of traffic management expertise. Furthermore, The integration of policy refinement corrects the reasoning trajectories and decisions of the backbone LLM, resulting in a discernible enhancement. This implementation attains SOTA or competitive performance across all the datasets and significantly surpasses GPT-4, accentuating the promising prospects for advancing intelligent transportation-oriented LLMs.
(a) The comparison on Jinan datasets.
(b) The comparison on Hangzhou datasets
Figure 4. The transferability of different methods.
Figure 5. The performance in the large-scale road network.
4.6. Generlization Comparision (RQ2)
4.6.1. Transferability
We first study the transferability ability of different methods by implementing the pre-trained model in another distinct road network. The experiment results are shown in Figure 4. Models labeled ”-T ” are pre-trained on a distinct road network (e.g., we evaluate the transferability in Hangzhou using the models pre-trained in Jinan). Otherwise, they are trained and tested in the same dataset. We observe a notable decline in the performance of reinforcement learning (RL)-based methods after the transfer, particularly evident in the Jinan 1 and Hangzhou 1 datasets. In contrast, LLMLight (LightGPT) consistently achieves superior performance and demonstrates remarkable transferability across all datasets. These findings highlight LightGPT’s exceptional generalization capability across diverse traffic contexts and robustness for practical implementations.
4.6.2. Scalability
Second, we delve into the scalability analysis of both traditional methods and LLMLight (LightGPT), scrutinizing their performance when applied to a considerably larger road network, New York, compared to the original road networks they are pre-trained on (i.e., Jinan and Hangzhou). The results of our experiments are detailed in Figure 5. We can observe significant performance degradations happening on most RL methods, as they demonstrate much worse performance than the heuristic method, Maxpressure. While the best RL methods exhibit suboptimal average travel time (ATT) in contrast to LLMLight (LightGPT), it is noteworthy that their policies result in significantly prolonged waiting times. Specifically, the waiting times are 57.80% and 49.80% longer than those achieved by ours. This observation suggests that the mechanism of the RL methods focuses on minimizing the overall number of queuing vehicles, possibly at the expense of causing extended waiting times for certain queues with fewer vehicles. The significance of minimizing waiting times cannot be overlooked in practical scenarios, as prolonged waits can induce anxiety and constitute a serious concern about the fairness of different drivers. In contrast, our proposed method ensures both the shortest travel and waiting times, highlighting its remarkable scalability and applicability when transitioning to much larger road networks.
4.6.3. Extreme High-traffic Scenarios
Next, we delve into an uncommon scenario when considerable traffic flow continuously appears at the intersection, a situation that rarely appears during training. To assess the efficacy of different methods in extreme high-traffic scenarios, we generate two synthetic traffic flow datasets on the Jinan and Hangzhou road networks, featuring an approximately fourfold increase in vehicles arriving within each 5-minute interval compared to the original datasets. Figure 6 shows the performance of both traditional methods and LLMLight armed with LightGPT (all the models are trained on Jinan 1 and Hangzhou 1 datasets, respectively). Similar to experiments on scalability, RL methods also showcase significant performance degradations, which exhibit poorer results than Maxpressure. This suggests that these pre-trained RL methods struggle to handle extreme high-traffic conditions when confronted with a significantly increased volume of vehicles compared to their training stage. In contrast, LLMLight (LightGPT) consistently shows superior performance, underscoring its robustness and practicality in much heavier traffic conditions.
Figure 6. The performance in extreme high-traffic scenarios.
Figure 7. The decision-making process of LightGPT.
4.7. Interpretability of LLMLight (RQ3)
To provide an in-depth assessment of LLMLight’s interpretability, we conducted a simulation case on the Hangzhou dataset. In Figure 7, we present the decision-making process of LightGPT. The detailed traffic conditions can be found in the Appendix A.3 (Table 3). In this simulated scenario, the north section experiences severe congestion, with a notable accumulation of queuing and approaching vehicles. LightGPT initiates its reasoning by comprehending the task and analyzing the traffic conditions at the targeted intersection. Subsequently, it adeptly identifies that the allowed lanes of signal NLSL demonstrate the heaviest congestion situation, thereby selecting it as the optimal signal phase to set. Unlike RL methods, LLMLight stands out not only for making effective control policies but also for its capacity to provide detailed explanations for each decision. This distinctive feature enhances the transparency and interpretability of LLMLight, contributing to a more informed understanding of its actions.
Figure 8. Prototype system.
4.8. Prototype System
We have successfully deployed a demonstration system designed to oversee the controlled traffic intersection and observe the decision-making of LightGPT. Figure 8 captures a snapshot of our demo system [^3] Our demo is available at https://llmtscs-demo.vercel.app/. In this setup, we provide users with real-time visualizations of the current traffic conditions on the right, along with statistical data of vehicles in the upper left corner. Additionally, the system further presents the reasoning behind LightGPT’s decisions through a message box in the interface’s lower left section.
5. Related Work
5.1. Traffic Signal Control
Traffic signal control (TSC) methods can be broadly categorized into transportation (Koonce and Rodegerdts, 2008; Hunt et al., 1982) and reinforcement learning (RL)-based approaches (Chen et al., 2020; Wu et al., 2023). Initially, FixedTime (Koonce and Rodegerdts, 2008) employs a fixed cycle length and phase split to regulate traffic signals in a predetermined pattern. Maxpressure (Varaiya, 2013) greedily selects signal phases to maximize road network throughput. PressLight (Wei et al., 2019a) inherits the concept of Maxpressure and combines it with a RL algorithm to optimize the throughput of the intersection. Additionally, some studies propose to design more representative state features to describe traffic conditions. Advanced-MP (Zhang et al., 2022b) studies vehicle movement-level state features, designing Advanced-MP features based on traffic demand.
5.2. Intelligent Transportation
The application of intelligent transportation mainly focuses on forecasting and control. Within traffic flow forecasting (Liu et al., 2023, 2022; Ning et al., 2023), (Li et al., 2018) advocates for using graph neural networks (GNNs) to facilitate the exchange of information among proximate regions. Regarding control tasks, NavTL (Sun et al., 2023) presents an RL-based framework to manage traffic signal activation and autonomous vehicle rerouting in hybrid traffic scenarios. (Xu et al., 2022) developed a Double-Deep Q-Network (DDQN)-based model with fine-grained state features for efficient on-ramp traffic flow control. To facilitate sustainable transportation, (Zhang et al., 2022a) develop dynamic charging pricing for electric vehicles (EV) via a multi-agent graph convolutional RL model. (Lai et al., 2023) propose a meta-learning-based framework to provide personalized vehicle energy consumption predictions for various drivers.
5.3. Large Language Models for Decision Making
LLMs recently emerged and exhibited a remarkable capacity for zero-shot reasoning and generalization. Exploiting this proficiency, numerous studies (Fu et al., 2023; Cui et al., 2023; Chen et al., 2023b; Wen et al., 2023; Jin et al., 2023; Schick et al., 2023; Sha et al., 2023; Wang et al., 2023b; Da et al., 2024; Xu et al., 2023a) have increasingly employed LLMs as central planners in diverse control tasks. For instance, GLAM (Carta et al., 2023) trains the LLM control agents through interactive engagement with the environment. In intelligent transportation, DriveGPT4 (Xu et al., 2023b) presents an end-to-end autonomous driving system capable of interpreting vehicle actions and addressing diverse questions posed by human users. More studies on utilizing LLMs in intelligent transportation can be found at (Zhou et al., 2023; Zhang et al., 2024).
6. Conclusion and Findings
In this study, we introduce LLMLight, a novel framework that leverages large language models (LLMs) as traffic signal control agents. By instructing LLMs to conduct a human-like, step-by-step analysis of current traffic conditions, the intelligent agent could judiciously choose the optimal control action, thereby enhancing the overall traffic efficiency at the intersection. Furthermore, we propose a backbone LLM training procedure for optimizing a traffic signal control-oriented LLM, LightGPT. Through extensive experiments conducted on nine traffic flow datasets, the framework armed with LightGPT demonstrates remarkable improvements compared to traditional methods, underlying its exceptional effectiveness and generalization ability across diverse traffic scenarios.
Our findings are summarized as follows: (1) Traditional traffic signal control methods often exhibit limited generalization capabilities, leading to suboptimal transferability, scalability, and effectiveness in unconventional scenarios. (2) While general LLMs often lack a deep understanding of traffic management, employing a tailored prompting workflow and domain-specific fine-tuning has proven effective in overcoming these limitations and improving the generated control policies. (3) The LLM-empowered traffic signal control not only demonstrates superior performance across diverse traffic scenarios but also can offer detailed explanations for each decision. This would significantly contribute to a more comprehensible and accountable traffic control system.
Appendix A Appendix
A.1. Visualizations of Intersection, Lanes, and Signal Phases
We present the most used setting of intersection, lanes, and signal phases in Figure 9.
A.2. Compared Models
A.2.1. Transportation Methods
  • Random: A baseline policy randomly switches signal phases with a fixed duration.
  • FixedTime (Koonce and Rodegerdts, 2008): A policy gives a pre-defined fixed cycle length with phase time, widely implemented in most traffic scenarios.
  • Maxpressure (Varaiya, 2013): The state-of-the-art heretics traffic signal control method in the transportation field, which greedily relieves vehicles on the lanes with the highest pressure (a variable derived from the difference between upstream and downstream queue lengths).
A.2.2. RL-based Methods
  • MPLight (Chen et al., 2020): A reinforcement learning (RL)-based method that utilizes pressure as both observation and reward. It employs FRAP, a network structure specifically designed to manage unbalanced traffic flow, as the foundational model.
  • AttendLight (Oroojlooy et al., 2020): Leveraging an attention mechanism, this model constructs observation features and predicts phase transition probability.
  • PressLight (Wei et al., 2019a): Leveraging the concept of Maxpressure with DRL, optimizing the pressure of the intersection.
  • CoLight (Wei et al., 2019b): This model employs a graph attention network (GAT) to enable communication among neighboring intersections, enhancing coordination and decision-making.
  • Efficient-CoLight (Wu et al., 2021): Building upon the CoLight model, it integrates efficient pressure as an observation, optimizing its decision-making capabilities.
  • Advanced-CoLight (Zhang et al., 2022b): Building upon the CoLight model, it incorporates advanced traffic state features, achieving the SOTA performance in traffic signal control tasks.
A.2.3. Large Language Models
  • Llama 2 (Touvron et al., 2023): This is a series of pre-trained and finely-tuned large language models (LLMs) developed by the AI group at Meta, ranging in scale from 7 billion to 70 billion parameters. They are adaptable for various natural language generation tasks. In our experiments, we use Llama-2-13B-chat-hf and Llama-2-70B-chat-hf in our experiments.
  • ChatGPT-3.5 (OpenAI, 2023b): Developed by OpenAI, ChatGPT-3.5 is a chatbot explicitly designed for engaging in conversations, answering queries, and assisting with diverse tasks. We adopt gpt-3.5-turbo-0613 API from OpenAI.
  • GPT-4 (OpenAI, 2023a): As the newest advancement in LLM by OpenAI, this SOTA foundation model is more reliable and creative compared to GPT-3.5. It excels in handling significantly more nuanced instructions and tasks. We utilize gpt-4-0613 API in our experiments.
Figure 9. An illustration of intersection, lanes, and signal phases. Lanes sharing a common color signify non-conflicting movements, and the passage is allowed in the corresponding signal phase. Signal phases include ETWT (go-through from east and west), ELWL (left-turn from east and west), NTST (go-through from north and south), and NLSL (left-turn from north and south).
Table 3. The statistics of the simulation case.
Signal Allowed Lanes Queue Length Approaching
ETWT East and west go-through 5 5
ELWL East and west left-turn 0 1
NTST North and south go-through 2 7
NLSL North and south left-turn 7 3
A.3. Failure of Other LLMs
To delve into the failure of other LLMs, we present the reasoning processes of Llama-2 and ChatGPT-3.5 in Figure 10. Table 3 shows the detailed traffic conditions. This scenario shows heavy congestion in the north and south sections, characterized by a notable accumulation of queuing vehicles in the left-turn lane and a considerable influx of approaching vehicles in the go-through lane. The Llama-2 models exhibit a common limitation in traffic management expertise and instruction-following (they disregard the provided commonsense knowledge). Although they understand the optimal traffic signal needs to address the most congested lanes of the intersection, Llama-2 opts to prioritize approaching vehicles, neglecting the more critical factors, queuing vehicles. On the other hand, ChatGPT-3.5 manifests a substantial logical flaw by asserting that releasing lanes with the lowest vehicle count is most effective in enhancing traffic conditions. This observation may suggest a pronounced hallucination issue (Ji et al., 2023; Huang et al., 2023) of ChatGPT-3.5 in the context of traffic signal control.
A.4. Visualizations of Road Networks
The visualizations of the road networks are shown in Figure 11.
A.5. Prompt Template
We present the prompt template of LLMLight in Table 4.
Figure 10. The reasoning of Llama-2 models and ChatGPT-3.5.
(a) Dongfeng, Jinan, China.
(b) Gudang, Hangzhou, China.
(c) Upper East Side, Manhattan, New York, US.
Figure 11. The road network systems of the Jinan, Hangzhou, and New York datasets.
Table 4. The prompt template.
Prompt Template
A traffic light regulates a four-section intersection with northern, southern, eastern, and western sections, each containing two lanes: one for through traffic and one for left-turns. Each lane is further divided into three segments. Segment 1 is the closest to the intersection. Segment 2 is in the middle. Segment 3 is the farthest. In a lane, there may be early queued vehicles and approaching vehicles traveling in different segments. Early queued vehicles have arrived at the intersection and await passage permission. Approaching vehicles will arrive at the intersection in the future.
The traffic light has 4 signal phases. Each signal relieves vehicles’ flow in the group of two specific lanes. The state of the intersection is listed below. It describes:
- The group of lanes relieving vehicles’ flow under each traffic light phase.
- The number of early queued vehicles of the allowed lanes of each signal.
- The number of approaching vehicles in different segments of the allowed lanes of each signal.
Signal: ETWT
Allowed lanes: Eastern and western through lanes
- Early queued: 1 (East), 4 (West), 5 (Total)
- Segment 1: 0 (East), 0 (West), 0 (Total)
- Segment 2: 0 (East), 2 (West), 2 (Total)
- Segment 3: 2 (East), 1 (West), 3 (Total)
Continued on next page...
Signal: NTST
Allowed lanes: Northern and southern through lanes
- Early queued: 2 (North), 0 (South), 2 (Total)
- Segment 1: 1 (North), 0 (South), 1 (Total)
- Segment 2: 1 (North), 0 (South), 1 (Total)
- Segment 3: 4 (North), 1 (South), 5 (Total)
Signal: ELWL
Allowed lanes: Eastern and western left-turn lanes
- Early queued: 0 (East), 0 (West), 0 (Total)
- Segment 1: 0 (East), 0 (West), 0 (Total)
- Segment 2: 0 (East), 0 (West), 0 (Total)
- Segment 3: 0 (East), 1 (West), 1 (Total)
Signal: NLSL
Allowed lanes: Northern and southern left-turn lanes
- Early queued: 4 (North), 3 (South), 7 (Total)
- Segment 1: 0 (North), 0 (South), 0 (Total)
- Segment 2: 0 (North), 0 (South), 0 (Total)
- Segment 3: 1 (North), 2 (South), 3 (Total)
Please answer:
Which is the most effective traffic signal that will most significantly improve the traffic condition during the next phase?
Note:
The traffic congestion is primarily dictated by the early queued vehicles, with the MOST significant impact. You MUST pay the MOST attention to lanes with long queue lengths. It is NOT URGENT to consider vehicles in distant segments since they are unlikely to reach the intersection soon.
Requirements:
- Let’s think step by step.
- You can only choose one of the signals listed above.
- You must follow the following steps to provide your analysis: Step 1: Provide your analysis for identifying the optimal traffic signal. Step 2: Answer your chosen signal.
- Your choice can only be given after finishing the analysis.
- Your choice must be identified by the tag: <signal>YOUR_CHOICE</signal>.